# Shanghai, China, 200241

### v2

@@ 3.2.1 $SU(2)_{0}$
where $\phi$ is the Coulomb parameter and $\mathfrak{q}$ is the instanton counting parameter. By inversing the mirror maps, we find the VEV for the Wilson loop is
@@ 4.2 Mordell-Weil group of special geometry and 2-group symmetry
The above conjectures are motivated by the conjecture in Closset:Uplane to which they reduce almost trivially when $r=1$. We point out that these conjectures can potentially be verified by studying carefully the charged spectrum as was done in rank-1 case in (92) since the root lattice of the flavor group has a nice embedding into $\mathscr{A}_{b}^{\vee}$. Furthermore, the 2-group symmetries of these higher rank cases can certainly be cross-checked against our B-model approach purposed in Section 3.1, since after all the complex structure moduli can be viewed as functions on at least locally on $\mathscr{B}$, e.g. they can be written explicitly as (36). Another interesting issue is to relate the Mordell-Weil torsion of $\mathscr{X}$ with the monodromy in the complex structure moduli space as heuristically outlined in Section 1.1 and further as is concretely exemplified by (36). We will leave a more detailed discussion of these issues in future works.

### v3

@@ 4.2.1 $SU(2)_{0}$
where $\phi$ is the Coulomb parameter and $\mathfrak{q}$ is the instanton counting parameter. By inverting the mirror maps, we find the VEV for the Wilson loop is
@@ 5.2 Mordell-Weil group of special geometry and 2-group symmetry
The above conjectures are motivated by the conjecture in Closset:Uplane to which they reduce almost trivially when $r=1$. We point out that these conjectures can potentially be verified by studying carefully the charged spectrum as was done in rank-1 case in (94) since the root lattice of the flavor group has a nice embedding into $\mathscr{A}_{b}^{\vee}$. Furthermore, the 2-group symmetries of these higher rank cases can certainly be cross-checked against our B-model approach purposed in Section 4.1, since after all the complex structure moduli can be viewed as functions on at least locally on $\mathscr{B}$, e.g. they can be written explicitly as (38). Another interesting issue is to relate the Mordell-Weil torsion of $\mathscr{X}$ with the monodromy in the complex structure moduli space as heuristically outlined in Section 2 and further as is concretely exemplified by (38). We will leave a more detailed discussion of these issues in future works.
@@ 6 Conclusion and discussion
One advantage of our discussion is that it can be readily applied to the study of 5D supergravity theories arising from M-theory compactifications on compact Calabi-Yau threefolds. For example, in a recent work Huang:2025xkc , the authors compute the refined BPS numbers of such threefolds and demonstrate that a one-form symmetry naturally emerges in the weak gravity limit due to B-field shifts. This symmetry, in turn, constrains the Wilson loop expansion of the refined BPS partition functions in 5D supergravity theories.